OpenStudy (anonymous):

Lucas is applying to both Stanford and Cal-poly. The probability that he gets accepted to Stanford is 0.4. The probability that he gets accepted to Cal-Poly is 0.3. The probability that he is accepted to both schools is 0.15. a) What is the probability that he is accepted to one of the two schools? b) What is the probability that he does not get accepted at either school?

OpenStudy (anonymous):

@ash2326 ?

OpenStudy (anonymous):

got this as my answer: (a) .7/.15 = (4.67) (b) .8/.15= (5.3) is that it?

OpenStudy (ash2326):

just a min. I'll check it

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

Probability of getting accepted in one=P(stan)+P(cal)-P(both)=0.4+0.3-0.15 Can you find that now?

OpenStudy (anonymous):

Is it .1-0.15?

OpenStudy (ash2326):

0.4+0.3-0.15=0.55

OpenStudy (anonymous):

nm, is it : .55

OpenStudy (ash2326):

yes

OpenStudy (anonymous):

ok cool:) now, b?

OpenStudy (ash2326):

just a min

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

sorry. I made a mistake in a, let me check again :(

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

For a Probability of getting accepted in one=P(stan)+P(cal)-2*P(both)=0.4+0.3-2*0.15 = 0.4

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

For b Probability of getting accepted in none=1-(P(stan)+P(cal)-P(both))=1-(0.4+0.3-0.15) = 0.45

OpenStudy (anonymous):

thank you!!!!!